OpenStudy (anonymous):

Choice A: 3 nights and one meal for $ 250 Choice B: 3 nights and 6 meals for $330 A motel offers the specials shown. What price is the motel charging per night?

OpenStudy (anonymous):

Well, there can be other charges, I would assume that any extra charges the value would be x or a variable

OpenStudy (anonymous):

So you would set it up as A = (250/3)*x B = (330/3)+6(x)

OpenStudy (anonymous):

I hope that this helps you.